# Approximating Unrelated Machine Weighted Completion Time Using Iterative Rounding and Computer Assisted Proofs

### v1

@@ 4 Analyzing Approximation Ratio with Computer Assistance
In this section, we prove the $1.36$-approximation ratio of our algorithm for a fixed machine $i\in M$, by comparing the right side of (5) and the expectation of the right side of (6) over all random choices of $\beta$. Specifically, we fix $j^{*}$, and compare the term after $(\sigma_{j^{*}}-\sigma_{j^{*}+1})$ in (5), and the counterpart in (6). Therefore, we only consider jobs in $[j^{*}]$ in this section. We restrict all the configurations on $i$ to be subsets of $[j^{*}]$. With a slight abuse of notations, we still use $y_{if}$ values to denote the masses of the configurations, after restricting jobs to $[j^{*}]$. That is, for any $f\subseteq[j^{*}]$, the new $y_{if}$ value equals the old $\sum_{f^{\prime}\cap[j^{*}]=f}y_{if^{\prime}}$. Now we set $\rho=2$ and our analysis in this section is tailored to this value of $\rho$. The main lemma we prove in this section is:
@@
Notice that all the definitions depend on $\beta$, and thus are random. In particular, $\ln\frac{L}{2}$ is also uniformly distributed in $[0,\ln 2)$. With the notations defined, we proceed to the proof of Lemma 4.1. Our goal is to define a constant $\alpha_{L}\in[1,1.5]$ for every $L\in[2,4)$ so that
$Q+\frac{1}{2}\cdot L^{2}-\frac{1}{2}\cdot S$ $\leq\alpha_{L}\cdot\frac{1}{2}\Big{(}Q+F\Big{)},\text{for every }\beta\in[0,\ln 2),\text{and}$ (9)
@@
The second inequality follows by defining $k^{\prime\prime}=k+k^{\prime}$. Also a job $j\in f$ is counted towards $v_{k}(f)$ if and only if $j\in J_{k^{\prime\prime}}$, and when it is counted, we have $q_{j}\cdot t=p_{j}$. Therefore, it remains to set $\alpha_{L}$ values so that (9) and (10) hold. If we set $\alpha_{L}=1.5$ for every $L$, then (9) holds even without the $\frac{1}{2}\cdot S$ term on the left-side, as $Q\leq F$ and $L^{2}\leq F$. So we can always set $\alpha_{L}\leq 1.5$.
Capturing $\alpha_{L}$ using a Mathematical Program
Till the end of the section, we fix the random variable $\beta$, and thus $L$. We define a size-configuration to be a multi-set $g\subseteq{\mathbb{R}}_{\geq 0}$ of positive reals with $|g|\leq n$. As the name suggests, a size-configuration is obtained from a configuration by replacing each job with its size. For every size configuration $g$, we define $v_{0}(g),v_{1}(g)$ and $v_{2}(g)$ to be the sum of elements of $g$ in $[1,2),[2,4)$ and $[4,8)$ respectively. This corresponds to the definition of $v_{k}(f)$ for configurations $f$; but we only need to define $v_{0}(g),v_{1}(g)$ and $v_{2}(g)$ for size-configurations $g$. Let $\alpha\in[1,1.5]$ be fixed; this will be our target value for $\alpha_{L}$. Now we introduce a program which aims to check if $\alpha_{L}=\alpha$ is satisfies (9). In the program, we have a variable $y^{\prime}_{g}$ for every size-configuration $g$, and three variables $V_{0},V_{1}$ and $V_{2}$. The program is defined by the objective (11), and constraints (12-15). [^2] Notice the program has uncountable many variables, but (12) and (13) will require there are countably many variables with non-zero values.
@@
$\sum_{g}y^{\prime}_{g}$ $=1$ (12) $y^{\prime}_{g}$ $\geq 0$ $\forall g$ (13) $\sum_{g}y^{\prime}_{g}\cdot(\sum g)$ $=L$ (14) $\sum_{g}y^{\prime}_{g}v_{k}(g)-V_{k}$ $=0$ $\forall k\in\{0,1,2\}$ (15)
@@
We used that $\sum_{a\in g}a^{2}\leq\big{(}\sum g\big{)}^{2}$ and $L=\sum_{g}(\sum g)$. So, for $\alpha=\frac{3}{2}-\frac{2}{L^{2}}$, the value of the sub-program is non-positive.
Analysis of Sub-Program (17).
@@
$\inf_{\mu,\lambda_{1},\lambda_{2}\in{\mathbb{R}}_{\geq 0}}\quad\sup_{y^{\prime},V_{1},V_{2}}\qquad$ $\sum_{f}y^{\prime}_{g}\Big{(}\big{(}1-\frac{\alpha}{2}\big{)}\sum_{a\in g}a^{2}-\frac{\alpha}{2}\big{(}\sum g\big{)}^{2}\Big{)}+\frac{1}{2}L^{2}-\frac{1}{2}\big{(}1+V_{1}^{2}+V_{2}^{2}\big{)}$
@@
Under the sup operator, $y^{\prime}$ is over all vectors such that $\sum_{g}y^{\prime}_{g}=1$ and $y^{\prime}_{g}\geq 0,\forall g$, and $V_{1},V_{2}\in{\mathbb{R}}$. Notice that we do not require (14) and (15) to hold. Also, we could allow $\mu,\lambda_{1}$ and $\lambda_{2}$ to be in ${\mathbb{R}}$. We chose the signs to be $+$, --, and -- before them, and restrict them in ${\mathbb{R}}_{\geq 0}$; this only increases the quantity. For any for fixed $\mu,\lambda_{1},\lambda_{2}\in{\mathbb{R}}_{\geq 0}$ and $y^{\prime}$, the objective (19) is maximized when $V_{1}=\lambda_{1}$ and $V_{2}=\lambda_{2}$. Also, for fixed $\mu,\lambda_{1},\lambda_{2},V_{1}$ and $V_{2}$, the objective is linear in $y^{\prime}$. So it is maximized when $y^{\prime}$ is a vertex point of the simplex. That is $y^{\prime}_{g}=1$ for some size-configuration $g$, and $y^{\prime}_{g^{\prime}}=0$ for every $g^{\prime}\neq g$. Therefore, (19) is equal to
@@
Then we redefine $v_{1}(g)$ and $v_{2}(g)$ as the total value of type-1 and type-2 elements in $g$ respectively. Now, the element $2\in g$ can contribute to either $v_{1}(g)$ or $v_{2}(g)$. This slight modification will not change the value of (20). By Lemma 4.3, we can restrict $g$ to satisfy the following property: the sum of the numbers in $g$, excluding the smallest one, is less than $L$. We say a number $a>0$ is flexible if $a\notin\{2,4,8\}$. We then prove
@@
Therefore, $(b,b^{\prime})$ is a convex combination of $(l,r^{\prime})$ and $(r,l^{\prime})$. Say $(b,b^{\prime})=\beta(l,r^{\prime})+(1-\beta)(r,l^{\prime})$ for some $\beta\in[0,1]$. By changing $\{b,b^{\prime}\}$ to $\{l,r^{\prime}\}$ and $\{r,l^{\prime}\}$ respectively, we obtain two multi-sets $g^{\prime}$ and $g^{\prime\prime}$. Moreover, we can make the type of $l$ and $r$ the same as that of $b$, and the type of $l^{\prime}$ and $r^{\prime}$ the same as that of $b^{\prime}$. $g^{\prime}$ and $g^{\prime\prime}$ may contain 00, but this is not an issue as removing 00 from the set does not change the objective. Then we have $\sum_{a\in g}a^{2}<\beta\sum_{a\in g^{\prime}}a^{2}+\beta\sum_{a\in g^{\prime\prime}}a^{2}$, $(\sum g)=(\sum g^{\prime})=(\sum g^{\prime\prime})$, $v_{1}(g)=\beta v_{1}(g^{\prime})+(1-\beta)v_{1}(g^{\prime\prime})$ and $v_{2}(g)=\beta v_{2}(g^{\prime})+(1-\beta)v_{2}(g^{\prime\prime})$. As $1-\frac{\alpha}{2}\geq 0$, one of $g$ and $g^{\prime}$ has a larger objective than $g$ in (20), for any $\mu,\lambda_{1},\lambda_{2}\in{\mathbb{R}}_{\geq 0}$. Therefore, we can disregard $g$. ∎

### v2

@@ 4 Analyzing Approximation Ratio with Computer Assistance
In this section, we prove the $1.36$-approximation ratio of our algorithm for a fixed machine $i\in M$, by comparing the right side of (5) and the expectation of the right side of (6) over all random choices of $\beta$. Specifically, we fix $j^{*}$, and compare the term after $(\sigma_{j^{*}}-\sigma_{j^{*}+1})$ in (5), and the counterpart in (6).
Therefore, we only consider jobs in $[j^{*}]$ in this section. We restrict all the configurations on $i$ to be subsets of $[j^{*}]$. With a slight abuse of notations, we still use $y_{if}$ values to denote the masses of the configurations, after restricting jobs to $[j^{*}]$. That is, for any $f\subseteq[j^{*}]$, the new $y_{if}$ value equals the old $\sum_{f^{\prime}\cap[j^{*}]=f}y_{if^{\prime}}$. Now we set $\rho=2$ and our analysis in this section is tailored to this value of $\rho$. The main lemma we prove in this section is:
@@ 4.1 Notations
Notice that all the definitions depend on $\beta$, and thus are random. In particular, $\ln\frac{L}{2}$ is also uniformly distributed in $[0,\ln 2)$.
With the notations defined, we proceed to the proof of Lemma 4.1. Our goal is to define a constant $\alpha_{L}\in[1,1.5]$ for every $L\in[2,4)$ so that
$Q+\frac{1}{2}\cdot L^{2}-\frac{1}{2}\cdot S$ $\leq\alpha_{L}\cdot\frac{1}{2}\Big{(}Q+F\Big{)},\text{for every }\beta\in[1,2),\text{and}$ (9)
@@
The second inequality follows by defining $k^{\prime\prime}=k+k^{\prime}$. Also a job $j\in f$ is counted towards $v_{k}(f)$ if and only if $j\in J_{k^{\prime\prime}}$, and when it is counted, we have $q_{j}\cdot t=p_{j}$.
Therefore, it remains to set $\alpha_{L}$ values so that (9) and (10) hold. If we set $\alpha_{L}=1.5$ for every $L$, then (9) holds even without the $\frac{1}{2}\cdot S$ term on the left-side, as $Q\leq F$ and $L^{2}\leq F$. So we can always set $\alpha_{L}\leq 1.5$; this recovers the $1.5$-approximation ratio without using the negative correlation.
4.2 Capturing $\alpha_{L}$ using a Mathematical Program
Till the end of the section, we fix the random variable $\beta$, and thus $L$. We define a size-configuration to be a multi-set $g\subseteq{\mathbb{R}}_{\geq 0}$ of positive reals with $|g|\leq n$. As the name suggests, a size-configuration is obtained from a configuration by replacing each job with its size. For every size configuration $g$, we define $v_{0}(g),v_{1}(g)$ and $v_{2}(g)$ to be the sum of elements of $g$ in $[1,2),[2,4)$ and $[4,8)$ respectively. This corresponds to the definition of $v_{k}(f)$ for configurations $f$; but we only need to define $v_{0}(g),v_{1}(g)$ and $v_{2}(g)$ for size-configurations $g$.
Let $\alpha\in[1,1.5]$ be fixed; this will be our target value for $\alpha_{L}$. Now we introduce a program which aims to check if $\alpha_{L}=\alpha$ is satisfies (9). In the program, we have a variable $y^{\prime}_{g}$ for every size-configuration $g$, and three variables $V_{0},V_{1}$ and $V_{2}$. The program is defined by the objective (11), and constraints (12-15). [^2] Notice the program has uncountable many variables, but (12) and (13) will require there are countably many variables with non-zero values.
@@ 4.2 Capturing $\alpha_{L}$ using a Mathematical Program
$\sum_{g}y^{\prime}_{g}$ $=1$ (12)
$y^{\prime}_{g}$ $\geq 0$ $\forall g$ (13)
$\sum_{g}y^{\prime}_{g}\cdot(\sum g)$ $=L$ (14)
$\sum_{g}y^{\prime}_{g}v_{k}(g)-V_{k}$ $=0$ $\forall k\in\{0,1,2\}$ (15)
@@
We used that $\sum_{a\in g}a^{2}\leq\big{(}\sum g\big{)}^{2}$ and $L=\sum_{g}y^{\prime}_{g}(\sum g)$. So, for $\alpha=\frac{3}{2}-\frac{2}{L^{2}}$, the value of the sub-program is non-positive.
4.4 Analysis of Sub-Program (17).
@@
$\inf_{\mu,\lambda_{1},\lambda_{2}\in{\mathbb{R}}_{\geq 0}}\quad\sup_{y^{\prime},V_{1},V_{2}}\qquad$ $\sum_{g}y^{\prime}_{g}\Big{(}\big{(}1-\frac{\alpha}{2}\big{)}\sum_{a\in g}a^{2}-\frac{\alpha}{2}\big{(}\sum g\big{)}^{2}\Big{)}+\frac{1}{2}L^{2}-\frac{1}{2}\big{(}1+V_{1}^{2}+V_{2}^{2}\big{)}$
@@
Under the sup operator, $y^{\prime}$ is over all vectors such that $\sum_{g}y^{\prime}_{g}=1$ and $y^{\prime}_{g}\geq 0,\forall g$, and $V_{1},V_{2}\in{\mathbb{R}}$. Notice that we do not require (14) and (15) to hold. Also, we could allow $\mu,\lambda_{1}$ and $\lambda_{2}$ to be in ${\mathbb{R}}$. We chose the signs to be $+$, --, and -- before them, and restrict them in ${\mathbb{R}}_{\geq 0}$; this only increases the quantity.
For any for fixed $\mu,\lambda_{1},\lambda_{2}\in{\mathbb{R}}_{\geq 0}$ and $y^{\prime}$, the objective (19) is maximized when $V_{1}=\lambda_{1}$ and $V_{2}=\lambda_{2}$. Also, for fixed $\mu,\lambda_{1},\lambda_{2},V_{1}$ and $V_{2}$, the objective is linear in $y^{\prime}$. So it is maximized when $y^{\prime}$ is a vertex point of the simplex. That is $y^{\prime}_{g}=1$ for some size-configuration $g$, and $y^{\prime}_{g^{\prime}}=0$ for every $g^{\prime}\neq g$. Therefore, (19) is equal to
@@
Then we redefine $v_{1}(g)$ and $v_{2}(g)$ as the total value of type-1 and type-2 elements in $g$ respectively. Now, the element $2\in g$ can contribute to either $v_{1}(g)$ or $v_{2}(g)$. This slight modification will not change the value of (20).
By Lemma 4.3, we can restrict $g$ to satisfy the following property: the sum of the numbers in $g$, excluding the smallest one, is less than $L$. We say a number $a>0$ is flexible if $a\notin\{2,4,8\}$. We then prove
@@
Therefore, $(b,b^{\prime})$ is a convex combination of $(l,r^{\prime})$ and $(r,l^{\prime})$. Say $(b,b^{\prime})=\gamma(l,r^{\prime})+(1-\gamma)(r,l^{\prime})$ for some $\gamma\in[0,1]$. By changing $\{b,b^{\prime}\}$ to $\{l,r^{\prime}\}$ and $\{r,l^{\prime}\}$ respectively, we obtain two multi-sets $g^{\prime}$ and $g^{\prime\prime}$. Moreover, we can make the type of $l$ and $r$ the same as that of $b$, and the type of $l^{\prime}$ and $r^{\prime}$ the same as that of $b^{\prime}$. $g^{\prime}$ and $g^{\prime\prime}$ may contain 00, but this is not an issue as removing 00 from the set does not change the objective.
Then we have $\sum_{a\in g}a^{2}<\gamma\sum_{a\in g^{\prime}}a^{2}+(1-\gamma)\sum_{a\in g^{\prime\prime}}a^{2}$, $(\sum g)=(\sum g^{\prime})=(\sum g^{\prime\prime})$, $v_{1}(g)=\gamma v_{1}(g^{\prime})+(1-\gamma)v_{1}(g^{\prime\prime})$ and $v_{2}(g)=\gamma v_{2}(g^{\prime})+(1-\gamma)v_{2}(g^{\prime\prime})$. As $1-\frac{\alpha}{2}\geq 0$, one of $g^{\prime}$ and $g^{\prime\prime}$ has a larger objective than $g$ in (20), for any $\mu,\lambda_{1},\lambda_{2}\in{\mathbb{R}}_{\geq 0}$. Therefore, we can disregard $g$. ∎